OpenStudy (wade123):

Find the derivative of f(x) = 6/x at x = -2. @sweetsunray

OpenStudy (anonymous):

same principle as before f(x)=6x^(-1) f'(x)= (-1) 6 x^(-1-1) = -6 x^(-2) = -6/x^2 plug -2 in for x, write down the fraction, can you simplify it?

OpenStudy (wade123):

it would be -3/2? @sweetsunray

OpenStudy (anonymous):

yes